At the 2012 State of the Union address, President Obama highlighted the rising cost of attending college. According to College Board, the cost to attend a college or university increased by 80% from 1994 to 2012. If the cost in 2012 is estimated at $18000, find the cost in 1994.
Answer by jim_thompson5910(35256)

Let x = cost in 1994 (in thousands of dollars)

So if it costs 18 thousand dollars now (in 2012) and it has "increased by 80% from 1994 to 2012", this means that


(Cost in 1994) + (80% of cost in 1994) = Cost in 2012

x + 80% of x = 18

x + 0.8x = 18

1.8x = 18

x = 18/1.8

x = 10

Remember that x is in thousands of dollars.

So the average cost to attend college in 1994 is $10,000

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------